Project

General

Profile

Actions

RunPlan20150115 Beamline Status

Beam stop
  • Was Will be serviced Thursday
    • Service revealed small hydraulic leak; it should be okay
  • Checked before xmas, was ok

DS

  • Operates fine at high current
  • Fridge trips at switch on
    • Needs expert to switch it on, but this should not be a problem (probably needs an expert to switch it on anyway)
  • DS test will be on Monday 12/01/2015
NC Magnets
  • Water flow issue with flow to magnet fixed 12/01/2015
    • JG agrees there is an issue, probably flow metre is reading incorrectly
    • Affects D1 magnet primarily, makes D1 power supply trip
    • May be that water flow is also lower than it used to be
Proton absorber
  • No issues
Target
  • Electrical cable changed during target change, something has happened, target crew will check next week
    • Maybe target will make an alarm as ISIS is dead (in concert with run control); make sure Pierrick is aware that there will be no ISIS current
Run settings
  • Add "check that datamover works" to detector and DAQ section
  • Pierrick will run magnets on Friday, we will need MOM keys at 14:00
Training
  • Henry requests mandatory PPS retraining for all
  • Note target training won't have dipping in beam so will need to "simulate" beam
Activation study
  • Need a pre-run to time the target in to ISIS (as the target system has been switched out)
  • BLOC training is possibly done at the same time; Steve Boyd to comment

Updated by Rogers, Chris almost 9 years ago ยท 7 revisions